i don't know how to do this kind of problem. (2/3)x + 7 = (3/7)
Answer by mathbird(71)

(2/3)x + 7 = (3/7)
(2/3)x=(3/7)-7
(2/3)x=-46/7
2x=-(46/7)*3
2x=-138/7
x=-69/7